The volume V of a given mass of a gas varies kirectly as the temperature T and inversely as the pressure P. If V=231cm^3 when T=42degrees and P=20 kg/cm^2, what is the volume when T=30degrees and P=15kg/cm^2?
Answer by fractalier(2101)

The equation that governs this is
V = kT/P
Now plug in to find k
231 = k(42)/20
k = 110
Now rewrite it with the new k and the new data to find the new volume...
V = 110T/P
V = 110(30)/15
V = 220 cm^3